# FIRE: A Failure-Adaptive Reinforcement Learning Framework for Edge Computing Migrations

### v2

@@ 9 Simulation Experiments
In this section, we provide numerical evaluation for our algorithms ImRE, ImDQL, and ImACRE, validating, and going beyond the results in Section 6. In particular, we show that we have solved the main research challenges identified in Section 1. We show that our failure-adaptive importance sampling based algorithms ImRE, ImDQL and ImACRE converged to optimality. We also show that they were able to learn a policy which mitigates the high cost of server failures, unlike the baselines in which importance sampling or backups were not used, given a tradeoff of incurring a higher cost during normal states due to backup storage and migration. Moreover, we can adapt to varied risk tolerances without re-training. We perform our experiments with the help of real world traces [52].
We compare our algorithms with the following baselines, over 10 runs. Each of these baselines, besides the greedy baseline, may be trained with deep Q-learning or actor critic (to correspond to our algorithm). We will compare by applying the converged policy of each algorithm in an online scenario where the rare events occur at their true rate.
RL with No Importance-Sampling (NIS): This simulator is a traditional Q-learning algorithm where there is no importance sampling of rare events. Rare events are simulated at their true rate, and backups are a possible action. Comparison to this baseline indicates the role of importance sampling in helping the learned policy adapt to rare events.
@@
RL without Rare Events Sampled (RES): This simulator is a Q-learning algorithm under the “normal scenario”, in which rare events are not modelled at all, during training. Likewise, backups are not considered as a possible action. Comparison to this baseline indicates the importance of including failures in the digital twin training environment.
@@
Figure 5: Multiple user service profile migrations: Comparison of our deep Q-learning algorithm ImDQL, with baselines QL-NIS, QL-WBA and QL-RES, and greedy algorithm, in an online scenario. ImDQL yields much lower costs in rare states, yielding lower costs on average despite a slight increase in cost for normal states.
@@
Figure 6: Shared service profile migrations: Comparison of our deep Q-learning algorithm ImDQL, with baselines NIS, RES, and greedy algorithm, in three different settings. For all settings, our algorithm achieves a lower cost at rare states, and on average, than the baselines.
@@
Single User Service Profile Migrations: In our experiments, we have 9 base stations (edge access points). The user is highly mobile across these coverage areas, and its mobility pattern follows a transition matrix. The ns-3 network simulator is used [52] to simulate dynamic channel network conditions, to obtain latency (delay) costs across different user-server location pairs. These latencies are in the range (2,10). As both the latency costs and migration costs are functions of the distance between user and access points, we set the migration cost $m_{ij}=d^{comm}_{i,j}+\varepsilon$, where $\varepsilon\in(-0.5,0.5)$. We set the cost of failures to be -500, representing the high cost in terms of user experience or safety, and the true failure rate as 0.01. The learning rate used is $0.05$ and both the actor and critic networks, for both our algorithms and the baselines, have a structure of $[24,48,24]$ nodes per layer.
@@
We apply our importance sampling based actor critic algorithm ImACRE and the corresponding actor critic baselines NIS, WBA, RES, for this scenario. In Fig. 3a, we show the convergence of ImACRE. In Fig. 4, we plot the overall average cost, average cost at rare states, and average cost at normal states, for our importance sampling based actor critic algorithm ImACRE and the corresponding actor critic baselines NIS, WBA, RES. All algorithms are run 10 times, and the average is plot. Our algorithm achieves a lower average cost at rare states, because ImACRE sufficiently samples rare events to learn a policy which prepares for them. Our algorithm incurs a trade-off in terms of the higher normal state cost, especially incurred through the migration and storage of backups. It can be seen in Fig. 4b that Algorithm NIS avoids the elevated costs associated with rare events in certain runs. However, its performance exhibits a greater variability. WBA and RES incur the highest costs in rare events due to the absence of backup usage and the fact that rare events are entirely excluded from sampling during simulator training in the case of RES.
@@
Multiple User Service Profile Migrations: In this scenario, we consider service profile migration for multiple users, where each user has their own service profile. We model 13 users mobile across 9 access points. Parameter settings are similar to the setup above. The natural failure rate is set to be 0.01. Users are mobile across the 9 APs, and their mobility patterns are sampled according to a transition probability matrix. We model different server failure types, corresponding to the differing amounts of maintenance resources available at each AP, which result in different amounts of time taken for the servers to come back online. For APs 1,6, and 8, we let their service downtime be 2 time-slots per failure. For all other APs, we let their service downtime be 1 time-slot per failure. We apply algorithm ImDQL, our importance sampling based deep Q learning algorithm in this scenario.
In Fig. 5, we see that our ImDQL policy outperforms the baselines by achieving a lower average cost during rare states, with a trade-off of a higher average cost in normal (non-rare) states. Overall, our algorithm still achieves a lower cost, when averaged over all timeslots. QL-NIS and the greedy algorithm perform better than the other baselines, as backups are still a potential action for the algorithm in QL-NIS, and the greedy algorithm places one backup at the second most likely location every time-slot. Nevertheless these algorithms do not avoid the rare event cost as much as our algorithm does.
@@
Shared service profiles: In this scenario (Fig. 6), we consider multiple users sharing the same service profile (SP). We model 100 users, with 70 users using SP Type 1, and 30 using SP Type 2. These users are spread over a 2-city map, which 2APs in one city and 3 in another. It has a failure cost of 1000 and failure rate of 1e-2. If a user’s required access point is in another city, the delay cost between cities will be nearly as large as the failure cost, which is much larger than the delay cost when accessing a service profile within the same city. We apply ImDQL to this scenario and compare its performance with NIS, RES and the greedy baseline. The greedy baseline deploys SPs to the $N$ most probable places w.r.t distribution of user transitions, where we search through all $N$ and plot the results with the lowest costs.
@@
We look at 3 settings: a) Full Deployment, where users are spread across zones evenly. The greedy baseline’s optimal solution is to fully deploy all SPs. b) Periodic, where users move across areas periodically, and c) Static, a corner case where users stay in their current zone. As seen in Fig. 6, in all settings, ImDQL achieves a lower cost in rare event states, followed by GRD (greedy), as compared to QL-NIS and QL-RES, which place little or no backups. While there is a trade-off with a higher cost in normal states, on average over all states, ImDQL and GRD still achieve a lower cost (Fig. 6 a) than the other baselines. ImDQL outperforms GRD in the periodic and static settings, as the greedy baseline cannot adapt to the presence of rare failures that may occur in each city.
@@
Catering to varying risk tolerances: We apply our algorithm RiTA (Algorithm 5), which caters to users having different risk tolerances. Given the risk tolerance level of a user $\zeta$, with probability $1-\zeta$, RiTA will select an action according to the converged policy of our algorithm ImRE (risk adverse component), and with probability $\zeta$ RiTA will select an action according to the converged policy of the baseline QL-RES (risk taking component). We apply RiTA to the single user (3 APs) service migration setting, which is a special case of the individual SP scenario, and to the multiple user shared SP scenario (5 APs across 2 cities).
@@
In Fig. 7 we present how the risk tolerance level $\zeta$ impacts the reward levels. As seen in Figs. 7 a) and b), for the single user individual SP scenario, the higher the risk tolerance, the higher the cost experienced at rare states (as it is less likely backups were placed). For normal states, as the risk tolerance level $\zeta$ decreases from $\zeta=1$, the cost first increases due to the increase in storage of backups (resulting from being increasingly risk adverse). After $\zeta=0.3$, the cost decreases. This is because as $\zeta$ decreases, the weightage given to the risk adverse component (our algorithm ImRE) increases. While ImRE incurs a higher storage cost than RES, it manages to learn a policy in which frequent migrations do not occur, hence decreasing the normal state cost. For the shared service profile scenario, as the risk tolerance $\zeta$ increases, the cost at normal state decreases due to the placement of less services, while the average cost (over all states) increases. This is because multiple users share service profiles like game environments, multiplying the cost when failures occur.
@@ 10 Conclusion
In edge computing, server failures may spontaneously occur, which complicates resource allocation and service migration. While failures are considered as rare events, they have an impact on the smooth and safe functioning of edge computing’s latency sensitive applications. As these failures occur at a low probability, it is difficult to jointly plan or learn an optimal service migration solution for both the typical and rare event scenarios. Therefore, we introduce a rare events adaptive resilience framework named FIRE, using the placement of backups and importance sampling based reinforcement learning, which alters the sampling rate of rare events in order to learn an optimal risk-averse policy. We prove the boundedness and convergence to optimality of our proposed tabular Q-learning algorithm ImRE. To handle large and combinatorial state and action spaces common in real-world networks, we propose deep Q-learning (ImDQL) and actor critic (ImACRE) versions of our algorithm. Furthermore, we propose a decision making algorithm (RiTA) that caters to heterogeneous risk tolerances across users. Finally, we use trace driven experiments to show that our algorithms converge to optimality, and are resilient towards server failures, unlike several baselines, subject to a trade-off in terms of a potential higher normal state cost. Our framework can also be modified for other resource allocation problems pertaining rare events with large consequences in communication and networking.

### v3

@@ 9 Simulation Experiments
In this section, we provide numerical evaluation for our algorithms FIRE-ImRE, FIRE-ImDQL, and FIRE-ImACRE, validating the results in Section 6. In particular, we show that we have solved the main research challenges identified in Section 1. We show that FIRE-ImRE, FIRE-ImDQL and FIRE-ImACRE converged to optimality. We also show that they were able to learn a policy which mitigates the high cost of server failures, unlike the baselines in which importance sampling or backups were not used, given a tradeoff of incurring a higher cost during normal states due to backup storage and migration. Moreover, we can adapt to varied risk tolerances without re-training. We perform our experiments with the help of real world traces [39], to obtain latency (delay) costs across different user-server location pairs in the network.
We compare our algorithms with the following baselines, over 10 runs. Each of these baselines, besides the greedy baseline, may be trained with deep Q-learning or actor critic (to correspond to our algorithm). We will compare by applying the converged policy of each algorithm in an online scenario where rare events occur at their true rate.
RL with No Importance-Sampling (NIS): This simulator is a traditional Q-learning algorithm, without importance sampling of rare events. Rare events are simulated at their true rate, and backups are a possible action. Comparison to this baseline indicates the role of importance sampling in helping the learned policy adapt to rare events.
@@
RL without Rare Events Sampled (RES): This simulator is a Q-learning algorithm under the “normal scenario”, where rare events and backups are not considered during training. Comparison to this baseline indicates the importance of including failures in the digital twin training environment.
@@
Figure 5: Multiple user service profile migrations: Comparison of our deep Q-learning algorithm FIRE-ImDQL, with baselines QL-NIS, QL-WBA and QL-RES, and greedy algorithm, in an online scenario. FIRE-ImDQL yields much lower costs in rare states, yielding lower costs on average despite a slight increase in cost for normal states.
@@
Figure 6: Shared service profile migrations: Comparison of our deep Q-learning algorithm FIRE-ImDQL, with baselines NIS, RES, and greedy algorithm, in three different settings. For all settings, our algorithm achieves a lower cost at rare states, and on average, than the baselines.
@@
Single User Service Profile Migrations: In our experiments, we have 9 base stations (edge access points). The user is highly mobile across these coverage areas, and its mobility pattern follows a transition matrix. The ns-3 network simulator is used [39] to simulate dynamic channel network conditions, to obtain latency (delay) costs across different user-server location pairs. These latencies are in the range (2,10). As both the latency costs and migration costs are functions of the distance between user and access points, we set the migration cost $m_{ij}=d^{comm}_{i,j}+\varepsilon$, where $\varepsilon\in(-0.5,0.5)$. We set the cost of failures to be -500, representing the high cost in terms of user experience or safety, and the true failure rate as 0.01 [41]. We choose this failure rate, because a higher order of 0.1 will entail that the failures are no longer rare events. And at a lower order of 0.001, the failures would give a smaller expected cost to the system. The learning rate used is $0.05$ and both the actor and critic networks, for both our algorithms and the baselines, have a structure of $[24,48,24]$ nodes per layer.
@@
We apply our importance sampling based actor critic algorithm FIRE-ImACRE and the corresponding actor critic baselines NIS, WBA, RES, for this scenario. In Fig. 3a, we show the convergence of FIRE-ImACRE. In Fig. 4, we plot the overall average cost, average cost at rare states, and at normal states, for our importance sampling based actor critic algorithm FIRE-ImACRE and the corresponding actor critic baselines NIS, WBA, RES. All algorithms are run 10 times, and the average is plot. Our algorithm achieves a lower average cost at rare states, because FIRE-ImACRE sufficiently samples rare events to learn a policy which prepares for them. Our algorithm incurs a trade-off with higher normal state costs, incurred through the migration and storage of backups. It can be seen in Fig. 4b that Algorithm NIS avoids the elevated costs associated with rare events in certain runs. However, its performance exhibits a greater variability. WBA and RES incur the highest costs in rare events due to the lack of backup usage and the exclusion of rare events from sampling during RES simulator training.
@@
Multiple User Service Profile Migrations: Here, we consider service profile migration for multiple users, each with their own service profile. We model 13 users mobile across 9 access points. Parameter settings are similar to the setup above. The natural failure rate is set to be 0.01. User mobility patterns are sampled according to a transition probability matrix. We model different server failure types, corresponding to the differing amounts of maintenance resources at each AP, which affect the time taken for the servers to come back online. For APs 1,6, and 8, we let their service downtime be 2 time-slots per failure. For all other APs, we let their service downtime be 1 time-slot per failure. We apply algorithm FIRE-ImDQL, our importance sampling based deep Q learning algorithm in this scenario. In Fig. 5, we see that our FIRE-ImDQL policy outperforms the baselines by achieving a lower average cost during rare states, with a trade-off of a higher average cost in normal (non-rare) states. Overall, our algorithm still achieves a lower cost, when averaged over all timeslots. QL-NIS and the greedy algorithm perform better than the other baselines, as backups are still a potential action for the algorithm in QL-NIS, and the greedy algorithm places one backup at the second most likely location every time-slot. Nevertheless these algorithms do not avoid the rare event cost as much as our algorithm does.
@@
Shared service profiles: In this scenario (Fig. 6), we consider multiple users sharing the same service profile (SP). We model 100 users, with 70 users using SP Type 1, and 30 using SP Type 2. These users are spread over a 2-city map, with 2APs in one city and 3 in another. It has a failure cost of 1000 and failure rate of 1e-2 [41]. If a user’s required access point is in another city, the delay cost between cities will be nearly as large as the failure cost, which is larger than the delay cost when accessing an SP within the same city. We apply FIRE-ImDQL to this scenario and compare its performance with NIS, RES and the greedy baseline. The greedy baseline deploys SPs to the $N$ most probable places w.r.t distribution of user transitions, where we search through all $N$ and plot the results with the lowest costs.
@@
We look at 3 settings: a) Full Deployment, where users are spread across zones evenly. The greedy baseline’s optimal solution is to fully deploy all SPs. b) Periodic, where users move across areas periodically, and c) Static, a corner case where users stay in their current zone. As seen in Fig. 6, in all settings, FIRE-ImDQL achieves a lower cost in rare event states, followed by GRD (greedy), as compared to QL-NIS and QL-RES, which place little or no backups. While there is a trade-off with a higher cost in normal states, on average over all states, FIRE-ImDQL and GRD still achieve a lower cost (Fig. 6 a) than the other baselines. FIRE-ImDQL outperforms GRD in the periodic and static settings, as the greedy baseline cannot adapt to the presence of rare failures that may occur in each city.
@@
Catering to varying risk tolerances: We apply our algorithm RiTA (Algorithm 4), which caters to users having different risk tolerances. Given the risk tolerance level of a user $\zeta$, with probability $1-\zeta$, RiTA will select an action according to the converged policy of our algorithm FIRE-ImRE (risk adverse component), and with probability $\zeta$ RiTA will select an action according to the converged policy of the baseline QL-RES (risk taking component). We apply RiTA to the single user (3 APs) service migration setting, which is a special case of the individual SP scenario, and to the multiple user shared SP scenario (5 APs across 2 cities).
@@
In Fig. 7 we present how the risk tolerance level $\zeta$ impacts the reward levels. As seen in Figs. 7 a) and b), for the single user individual SP scenario, the higher the risk tolerance, the higher the cost experienced at rare states (as it is less likely backups were placed). For normal states, as the risk tolerance level $\zeta$ decreases from $\zeta=1$, the cost first increases due to the increase in storage of backups (resulting from being increasingly risk adverse). After $\zeta=0.3$, the cost decreases. This is because as $\zeta$ decreases, the weightage given to the risk adverse component (our algorithm FIRE-ImRE) increases. While FIRE-ImRE incurs a higher storage cost than RES, it manages to learn a policy in which frequent migrations do not occur, hence decreasing the normal state cost. For the shared service profile scenario, as the risk tolerance $\zeta$ increases, the cost at normal state decreases due to the placement of less services, while the average cost (over all states) increases. This is because multiple users share service profiles like game environments, multiplying the cost when failures occur.
@@ 10 Conclusion
In edge computing, server failures may spontaneously occur, which complicates resource allocation and service migration. While failures are considered as rare events, they have an impact on the smooth and safe functioning of edge computing’s latency sensitive applications. As these failures occur at a low probability, it is difficult to jointly plan or learn an optimal service migration solution for both the typical and rare event scenarios. Therefore, we introduce a rare events adaptive resilience framework named FIRE, using the placement of backups and importance sampling based reinforcement learning, which alters the sampling rate of rare events in order to learn an optimal risk-averse policy. We prove the boundedness and convergence to optimality of our proposed tabular Q-learning algorithm FIRE-ImRE. To handle large and combinatorial state and action spaces common in real-world networks, we propose deep Q-learning (FIRE-ImDQL) and actor critic (FIRE-ImACRE) versions of our algorithm. Furthermore, we propose a decision making algorithm (RiTA) that caters to heterogeneous risk tolerances across users. Finally, we use trace driven experiments to show that our algorithms converge to optimality, and are resilient towards server failures, unlike several baselines, subject to a trade-off in terms of a potential higher normal state cost. Our framework can also be modified for other resource allocation problems pertaining rare events with large consequences in communication and networking.